# Rethinking Whole-Body CT Image Interpretation: An Abnormality-Centric Approach

### v1

@@ 5.2 Experiment Results
Model DSC SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
@@
MedULS + LLaVA-Med 14.53 6.10 9.54 2.18 0.19 27.35 82.98 11.45 11.48 9.20 1.61
LiSA + LLaVA-Med 17.48 21.35 10.57 2.41 0.28 29.18 83.47 11.59 12.34 9.80 1.61
BiomedParse + LLaVA-Med 16.57 19.87 10.66 2.38 0.26 29.66 83.39 11.74 12.62 9.86 2.16
OminiAbnorm-CT 32.40 58.43 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
@@
MedULS + LLaVA-Med 12.26 5.37 10.73 2.60 0.26 29.34 82.89 11.40 12.79 9.82 1.61
LiSA + LLaVA-Med 12.04 13.76 11.58 2.71 0.22 30.76 83.21 11.36 13.40 10.23 1.83
BiomedParse + LLaVA-Med 13.60 17.79 11.21 2.64 0.25 30.24 83.33 11.47 13.50 10.31 2.34
OminiAbnorm-CT 27.74 45.16 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
@@
MedULS + LLaVA-Med 11.89 7.66 10.25 2.69 0.29 27.40 82.49 11.24 12.17 9.53 1.23
LiSA + LLaVA-Med 11.23 16.67 11.14 2.72 0.30 29.79 82.88 11.15 13.36 10.07 1.42
BiomedParse + LLaVA-Med 15.35 17.57 11.07 2.72 0.39 31.12 83.01 11.66 13.85 10.40 1.78
OminiAbnorm-CT 29.89 45.41 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
Text-guided Grounded Report Generation. As shown in Table 3, OminiAbnorm-CT demonstrates superior grounding capabilities, with DSC scores improving by 14.24, 14.14, and 14.54 in axial, coronal, and sagittal planes respectively compared to the best baselines. For report generation, OminiAbnorm-CT excels across both lexical overlap metrics (average improvements of 1.3 on BLEU-1 and 2.2 on ROUGE-1 across all planes) and clinical relevance metrics (average gains of 4.0 on RaTEScore and 2.9 on RadGraph across all planes), confirming its superior ability to generate reports that align with clinical assessments. Additionally, OminiAbnorm-CT achieves significantly higher Specificity, demonstrating its robustness to queries about non-existing abnormalities.
5.3 Ablation Study
The integration of the segmentation module allows OminiAbnorm-CT to acquire grounding evidence for abnormal findings and generate descriptions accordingly. To validate its necessity, we conduct an ablation study comparing the full model against a variant where the segmentation module is removed (denoted as OminiAbnorm-CT w/o Seg). This modified version directly generates findings descriptions without anatomical localization guidance.
In the text-guided grounded report generation task, as shown in Table 4, OminiAbnorm-CT without the segmentation module demonstrates notable performance degradation, with inferior results on 23 of 27 metrics across three planes. Furthermore, detailed analysis reveals that without the segmentation module, the model struggles to accurately detect the queried abnormalities in CT images, exhibiting a pronounced negative bias that leads to high specificity scores.
Similarly, in the grounded report generation task, as depicted in Table 5, removing the segmentation module leads to consistent performance degradation on all the metrics, further confirming that segmentation-based evidence is crucial for generating high-quality medical reports.
Table 4: Ablation study of the segmentation module for text-guided grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
Model SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
@@ 5.3 Ablation Study
Table 5: Ablation study of the segmentation module in grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
OminiAbnorm-CT w/o Seg 7.73 4.74 3.16 27.87 73.14 17.65 15.18 13.58 3.42
OminiAbnorm-CT 19.00 12.26 8.40 33.85 86.24 22.45 22.27 19.43 4.60
Coronal (n=750)
OminiAbnorm-CT w/o Seg 8.12 4.81 3.12 31.03 75.55 15.64 15.27 13.23 3.84
OminiAbnorm-CT 19.28 12.03 7.96 35.40 85.56 20.81 22.35 18.47 4.65
Sagittal (n=591)
OminiAbnorm-CT w/o Seg 8.79 5.21 3.33 31.21 77.41 16.18 15.86 13.87 3.91
OminiAbnorm-CT 16.69 10.41 6.98 33.45 84.94 18.77 20.67 17.61 3.83
6 Related Work
Medical image segmentation aims to delineate clinically meaningful regions on medical images. For radiology images, most datasets focus on annotations for important organs and anatomical structures Totalsegmentator ; abdomenatlas ; FLARE ; AMOS22 or lesions heller2021kidney ; MSD ; BraTS2023GLI ; de2024uls23 ; yan2018deep . Specialist segmentation models nnUNET ; UNet ; zhou2023nnformer ; zhou2019unet++ have achieved remarkable success in the past decade, while recently, building foundational generalist models MedSAM ; zhao2024foundation ; zhao2023one ; du2024segvol has received increased attention. However, few works contribute to the grounding of abnormal image findings, which encompass a broader spectrum including abnormal organs, potential lesions, and more generally, any anomalies meaningful for clinical decision-making.
Medical image report generation aims to faithfully interpret the medical images with formal text, which is indispensable for numerous clinic procedures rao2025multimodal . Existing datasets with image-report pairs are limited to specific modalities and regions, such as chest X-ray johnson2019mimic , brain MRI lei2024autorg and chest CT hamamci2024foundation ; chen2024bimcv , overlooking the clinical demands for multi-plane whole-body CT image report generation. Recent progress zhang2024radgenome ; chen2025large ; lei2024autorg ; chen2024vision on grounded report generation are limited to structured organ-level segmentation and description, failing to localize and interprete fine-grained abnormalities.
Generative visual-language foundation models for medicine are built on large-scale multi-modal medical data, and have demonstrated exceptional performance and generalization capabilities across diverse medical tasks wu2023towards ; moor2023med ; zhang2024generalist ; li2023llava ; meddr . Despite their potential application to multi-plane, whole-body CT images, however, most existing approaches fail to simultaneously provide grounded evidence during generation, leading to poor explainability and hallucinations that are more difficult to detect.
7 Conclusion
This paper advances the automatic grounded interpretation of CT imaging from an abnormality-centric view. To support this, we develop a hierarchical taxonomy of 404 representative abnormal findings on CT images. We contribute OminiAbnorm-CT-14K, a meticulously annotated dataset with detailed grounding and description annotation for around 19K abnormalities, from 14.5K multi-plane CT images across the entire human body. Built on this dataset, OminiAbnorm-CT enables automatic grounding and description of abnormalities, while supporting flexible clinical usage with text queries or visual prompts. Comprehensive evaluations across three clinically representative tasks demonstrate OminiAbnorm-CTs superior performance. Our work pioneers abnormality-centric CT interpretation, enhancing diagnostic transparency through accurate grounding and detailed characterization of findings, with significant potential to transform radiological practice.
Acknowledgments and Disclosure of Funding
Weidi would like to acknowledge the funding from Scientific Research Innovation Capability Support Project for Young Faculty (ZY-GXQNJSKYCXNLZCXM-I22).
Appendix A Appendix
A.1 Qualitative Experiment Results
Fig. 4 presents several examples from the grounded report generation task, comparing the segmentation and report generation results between OminiAbnorm-CT and BiomedParse+LLaVA-Med. The results clearly demonstrate that BiomedParse fails to detect any abnormalities in the latter two cases, consequently leading LLaVA-Med to generate reports that are hardly relevant to the images. In contrast, OminiAbnorm-CT successfully localizes the abnormal regions across all examples and produces comparatively more accurate reports for the abnormalities.
In Fig. 5, we demonstrate several examples from the text-guided grounded report generation task, and compare the segmentation and report generation results of OminiAbnorm-CT and BiomedParse+LLaVA-Med. As illustrated in the first two examples, BiomedParse fails to accurately localize the queried abnormalities. While in the latter two examples, LLaVA-Med is unable to correctly interpret the segmentation results when generating reports. In contrast, OminiAbnorm-CT can successfully localize the queried abnormalities across all examples and generates more precise and clinically relevant reports.
In Fig. 6, we demonstrate several representative examples in the visual prompted report generation task, comparing the abnormality descriptions generated by OminiAbnorm-CT, LLaVA-Med, and QWen2.5-VL-7B. It shows that LLaVA-Med and QWen2.5-VL-7B exhibit some fundamental errors in their results. For instance, in the first case, LLaVA-Med misidentifies the kidney as the gallbladder; While in the last case, QWen2.5-VL-7B fails to distinguish between the liver and gallbladder. By contrast, OminiAbnorm-CT is able to precisely capture and identify the abnormal findings highlighted by the visual prompts and generate high quality reports.
A.2 Limitations
As a pioneering attempt toward automatic grounding and interpretation of abnormalities on multi-plane and whole-body CT images, our work certainly has several limitations: (i) Regarding data annotation, despite providing a large-scale dataset, some rare abnormalities (55 out of 404 in the taxonomy) remain uncovered, as shown in Section A.11. Additionally, many abnormalities have very few annotations, which significantly restricts the model’s capability range. Expanding annotations for these rare categories is treated as our future work. Meanwhile, the CT images in OminiAbnorm-CT-14K may also exhibit demographic and regional bias. (ii) Due to resource constraints, we currently provide grounding annotation for a few representative slices on each CT image, rather than full volumetric segmentation. In the future, we could potentially address this limitation by using interactive segmentation models such as nnInteractive [27]. (iii) Limited by the 3D image perception capability of multi-modal language models, which is also a common challenge in current generative visual-language foundation models for medicine [53, 44, 58, 33, 20], we only considered single slice as input when generating reports. Enhancing models’ ability to reason across adjacent slices and integrate volumetric context represents a critical direction for our future research.
A.3 More Implementation Details of OminiAbnorm-CT
Training Hyperparameters. We adopt 8 NVIDIA A100-80G GPUs for training, with 1 sample per device and gradient accumulation set to 4. We train OminiAbnorm-CT for 250K iterations in total, and use 5% steps for warming up. We take AdamW [36] as optimizer with a learning rate of 2e-4. When optimizing the multi-modal language model and the segmentation module jointly, we assign equal weights to the text generation loss and segmentation loss, as defined in Equation 7. We set the rank of LORA to 32.
Training Data. We pad and rescale all images to 512$\times$512. For training data sampling from four tasks: visual prompted report generation, grounded report generation, text-guided grounded report generation, and general VQA, we control the ratio to be 1:1:1:1. Additionally, in the abnormality grounding task, we maintain a ratio of 6:4 between OminiAbnorm-CT-14K and public lesion segmentation data. For each abnormality description, we use the following prompt to have GPT rewrite three different versions, then randomly select one as the ground truth:
A.4 Implementation Details of Baselines
In visual prompted generation task, we prompt each abnormality with all types of visual prompts, and take the maximum score. All the methods are prompted with the following template:
In grounded report generation task, we integrate LLaVA-Med with 2D segmentation models as baselines. Specifically, we re-implement the MedULS [13] with a 2D nnU-Net [26] and the public lesion segmentation datasets in Table 6, covering all the datasets in the official ULS-23 challenge and 10 additional ones. To our knowledge, this represents the segmentation model with the broadest capability range (in terms of abnormality variety) that can be constructed from currently available public datasets. For BiomedParse [61] and LiSA [31], we prompt them with ‘Abnormal findings on the CT image’ to derive segmentation results for all the abnormalities on the input CT image. Then, we combine each segmentation model with LLaVA-Med by converting the segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on these visual cues:
In text-guided grounded report generation task, we use the simulated text queries as prompts for BiomedParse and LiSA to derive segmentation results for the queried abnormalities, which are detailed in Section 3.4. For MedULS, since it doesn’t support text-prompted segmentation, we simply use its unconditioned segmentation results. Similarly, we convert their segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on them:
A.5 Quality Verification on OminiAbnorm-CT-14K
We conducted rigorous and comprehensive quality verification on the annotations in OminiAbnorm-CT-14K. For each annotator, we randomly sampled 100 annotated images on the axial plane, 50 on the coronal, and 50 on the sagittal plane. Then a senior radiologist with 12 years of experience assesses the annotation quality on four key metrics: (i) Detection rate measures the percentage of abnormalities successfully identified and annotated by the annotator, without any omission; (ii) Grounding precision evaluates the percentage of grounding annotations that properly encompass the primary regions of the abnormality, while minimizing false positive areas; (iii) Report concordance quantifies the percentage of description annotations that faithfully reflect the linked abnormal findings on the image; (iv) Classification accuracy measures the percentage of abnormalities that are correctly categorized.
Our verification results demonstrated that: (i) 3 out of 4 annotators achieved perfect scores across all four metrics (100%). (ii) Only one annotator had 10 abnormality annotations where the descriptions were not sufficiently accurate and the category labels were also incorrect, resulting in Report concordance and Classification accuracy of 95%, while maintaining 100% Detection rate and Grounding precision. These results confirm the high quality of annotations in our OminiAbnorm-CT-14K.
A.6 The Long-Tail Distribution of OminiAbnorm-CT-14K
To mitigate the long-tail distribution in OminiAbnorm-CT-14K, we identified underrepresented organs in our annotated corpus and strategically employed GPT-4o to analyze unlabeled reports. Using the following prompt, we efficiently filtered cases containing abnormal findings related to these underrepresented organs:
A.7 Comparison between OminiAbnorm-CT-14K and Existing Datasets
We compare OminiAbnorm-CT-14K with existing public CT image datasets in Table 6, including those widely used for lesion segmentation, lesion detection, organ segmentation, or report generation. In contrast to the substantial limitations exhibited by these datasets, as detailed in Sec. 3.1, OminiAbnorm-CT-14Kepresents the first large-scale dataset designed for abnormality grounding and description across multi-plane and whole-body CT images.
Table 6: Comparison of key characteristics between OminiAbnorm-CT-14K and widely-used public CT imaging datasets. Note that CCC18 and DeepLesion has no category label for each annotated lesion. Even though some scans in these datasets are isotropic, they are acquired as axial-plane. Abbreviations: Rpt = Report; Cat = Abnormality Category; Img = Image; Seg. = Segmentation; Det. = Detection; Gen = Generation.
Dataset Task Anatomy Plane Rpt #Cat #Img
[…]
Lesion Seg. Axial 340 10K
OminiAbnorm-CT-14K Lesion Det. Whole Body Coronal 255 2K
Report Gen. Sagittal 223 1.6K
A.8 Ablation Study on Context Range
In clinical practice, the contextual information provided by adjacent slices is critical for the interpretation of CT images, which is also widely validated in tasks such as medical image segmentation [26] and diagnosis [64]. To investigate the impact of context range on report generation, in the visual prompted report generation task, we adjust the number of input slices and evaluate its impact on a base QWen2.5-VL-7B model. As shown in Table 7, when we increase the input slices, there is no consistent performance improvement on any metric. We speculate this is mainly because multi-modal language models, which generally encode and project each input image into tokens independently, can hardly capture the subtle spatial relationships across adjacent slices. Therefore, we only take the center slice as image input to the multi-modal language model in the subsequent training and evaluation of OminiAbnorm-CT.
Table 7: Impact of multi-slice context range on visual prompted report generation, with the base Qwen2.5-VL-7B model as baseline. Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
Input Slices B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG

### v2

@@ 6.1 Experiment Settings
6.2 Evaluation Metrics
Existing Metrics. For grounding results, we calculate Dice Similarity Coefficient (DSC). For generated report, we include natural language generation metrics: BLEU (B) [bleu], BERTScore (BERTSc) [zhangbertscore], METEOR (MTR) [banerjee2005meteor], and ROUGE (R) [rouge]; as well as medical-specific metrics: RaTEScore (RaTESc) [zhao2024ratescore] and RadGraph (RadG) [delbrouck2024radgraph].
AbnormRubric (Rubric) is proposed as a metric for generated report. Unlike above metrics that are limited to simple word matching or unexplainable semantic similarity, AbnormRubric aligns with clinical judgment by comprehensively evaluating the abnormalities in the generated report from: (i) the detection ratio, (ii) the description accuracy of radiological attributions, and (iii) the hallucination. Formally, given all the ground-truth abnormalities ($\text{TP}+\text{FN}$), a LLM is first instructed to numerate the detected (TP) and hallucinated abnormalities (FP) in the generated report:
$\text{Recall}=\text{TP}/(\text{TP}+\text{FN})$ (9)
$\text{Precision}=\text{TP}/(\text{TP}+\text{FP})$ (10)
Then, for each detected abnormality, the LLM is instructed to judge whether the generated description align with ground-truth report on different radiological attributions:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\text{Attribution}_{i}\text{ aligns with GT}]$ (11)
where $\mathbb{I}$ is the indicator function. Notably, for each type of abnormality in our defined taxonomy, the radiologists have tailored and identified a set of radiological attributions, including location, size, morphology, and density, that are clinically important and should be accurately described. And the final score is formulated as:
$\text{Recall}^{*}=(1+\text{Accuracy})/2\times\text{Recall}$ (12)
$\text{Rubric}=2\times\frac{\text{Recall}^{*}\times\text{Precision}}{\text{Recall}^{*}+\text{Precision}}$ (13)
More details of AbnormRubric are in the Appendix.
6.3 Internal Validation Results
Visual Prompted Report Generation. Table 1 reports that OmniAbnorm-CT outperforms all baselines on 29 out of 30 metrics. Specifically, compared to the strongest baseline, OmniAbnorm-CT achieves average improvements of 2.4 in BLEU-1, 2.8 in RaTEScore and 4.26 in AbnormRubric across all three planes. It demonstrates OmniAbnorm-CT’s effectiveness in semi-automated workflows, where radiologists mark the abnormalities and the model provides accurate and detailed interpretations.
Grounded Report Generation. As shown in Table 2, OmniAbnorm-CT demonstrates superior grounding capabilities, outperforming the best baselines by 15.42, 14.04, and 19.56 in DSC on axial, coronal, and sagittal planes respectively. For report generation, OmniAbnorm-CT achieves the best performance across all metrics, with notable improvements in AbnormRubric, RadGraph and RaTEScore across all anatomical planes, indicating that the generated reports can better align with clinical standards. These results demonstrate OmniAbnorm-CT’s potential for fully automated abnormality localization and report generation in clinical settings.
Text-guided Grounded Report Generation. In Table 3, OmniAbnorm-CT achieves substantially better grounding results, improving DSC over the strongest baseline on axial (+14.24), coronal (+14.14), and sagittal (+14.54). For report generation, OmniAbnorm-CT outperforms in both general overlap metrics (average +1.3 BLEU-1, +2.2 ROUGE-1 across planes) and clinical metrics (average +14.42 AbnormRubric across planes), indicating closer alignment with clinical assessments. It indicate that OmniAbnorm-CT can effectively follows radiologist instructions to ground and interpret the relevant abnormalities.
Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
Visual prompted report generation (n=271)
GPT4o - 19.15 17.37 7.59 1.81 37.86 85.52 20.46 19.16 15.18 3.73
QWen25VL7B(OG) - 15.15 6.41 2.57 0.42 37.04 82.84 14.62 9.52 7.37 2.51
ViP-LLaVA - 10.31 17.59 6.20 0.55 32.88 85.46 15.33 19.18 16.36 1.95
MedDr ✓ - 13.52 3.24 1.13 0.17 20.36 54.28 3.58 4.94 4.29 0.74
LLaVA-Med ✓ - 12.22 17.63 7.40 1.20 34.96 85.52 19.81 19.01 15.64 2.92
BiomedGPT ✓ - 12.07 10.99 4.57 0.60 35.14 82.62 17.30 13.92 11.37 2.67
OmniAbnorm-CT ✓ - 24.09 19.09 7.71 1.67 39.87 86.64 19.46 21.02 17.34 4.27
Grounded report generation (n=271)
MedULS + LLaVA-Med 11.63 3.96 12.86 7.09 1.32 33.53 81.88 17.31 15.88 12.87 2.50
LiSA + LLaVA-Med 18.34 4.37 12.33 6.54 0.62 35.43 83.57 17.03 15.87 12.71 2.69
BiomedParse + LLaVA-Med 11.78 5.32 11.98 6.51 0.74 34.76 83.47 16.59 16.10 13.08 2.87
OmniAbnorm-CT ✓ 22.20 8.62 18.55 12.10 8.14 34.93 86.09 21.61 22.36 18.32 2.57
Text-guided grounded report generation (n=271)
MedULS + LLaVA-Med 11.69 3.23 9.47 2.86 0.33 25.57 74.55 11.57 10.81 8.91 1.32
LiSA + LLaVA-Med 9.72 0.17 2.85 1.43 0.89 4.68 20.26 3.08 3.32 2.90 1.40
BiomedParse + LLaVA-Med 10.50 0.33 4.33 2.06 1.44 6.65 25.52 4.50 4.77 4.48 1.63
OmniAbnorm-CT ✓ 14.13 15.12 12.38 3.76 1.18 32.25 76.34 13.55 14.85 11.97 3.55
Table 4: External validation results across three tasks. Metrics are averaged within each abnormality category and then across all categories. Within each block, the best result is bolded and the second best is underlined. For metrics that are not applicable to a given task, we report “-”. Models optimized with medical data are marked with ✓.
6.4 External Validation Results
As demonstrated in Table 4, on the external validation set, OmniAbnorm-CT maintains strong performance across all three tasks. In visual prompted report generation, OmniAbnorm-CT achieves the highest AbnormRubric score (+4.94 over the best baseline), indicating that its generated reports align more closely with clinical criteria. Moreover, in both grounded report generation and text-guided grounded report generation, OmniAbnorm-CT surpasses all baselines on report generation metrics and shows superior grounding ability (DSC improvements of +10.57 for grounded report generation and +2.44 for text-guided grounded report generation). These results suggest that OmniAbnorm-CT generalizes well to external datasets.
7 Conclusion
This paper advances the automatic grounded interpretation of CT imaging from an abnormality-centric view. To support this, we develop a hierarchical taxonomy of 404 representative abnormal findings. We contribute OmniAbnorm-CT-14K, a meticulously annotated dataset with detailed grounding and description annotation for around 19K abnormalities, from 14.5K multi-plane CT images across the entire human body. Built on this dataset, OmniAbnorm-CT enables automatic grounding and description of abnormalities, while supporting flexible clinical usage with text queries or visual prompts. We establish three representative tasks and devise a clinically grounded metric to demonstrate OmniAbnorm-CTs superior performance.
Limitations. As an early attempt, this paper certainly has limitations and future works to continue: On dataset, we annotate representative slices rather than full volumes, primarily due to the high cost of expert labeling. We will explore semi-automated tools to scale to full-volume annotations in future. For OmniAbnorm-CT, current VLM has limited volumetric perception, constrained by training data and compute resources. We plan to curate larger-scale whole-body CT datasets for pretraining and adapting a 3D vision encoder in future. Nevertheless, our work pioneers abnormality-centric CT interpretation, improving diagnostic transparency via accurate grounding and detailed characterization of abnormalities. We believe this paradigm shift holds substantial potential to transform radiology practice.
Rethinking Whole-Body CT Image Interpretation:
An Abnormality-Centric Approach
Appendix
Appendix A Qualitative Experiment Results
Fig. 4 presents several cases from the grounded report generation task, comparing the segmentation and report generation results between OmniAbnorm-CT and BiomedParse+LLaVA-Med. It clearly shows that BiomedParse fails to detect any abnormalities in the latter two cases, and consequently, LLaVA-Med generates reports irrelevant to the abnormalities. In contrast, OmniAbnorm-CT successfully localizes the abnormalities across all cases and produces more accurate reports.
In Fig. 5, we compare OmniAbnorm-CT with BiomedParse+LLaVA-Med on the text-guided grounded report generation task. BiomedParse mislocalizes the queried findings in first two cases, while LLaVA-Med fails to correctly interpret the segmentation results in last two cases. In contrast, OmniAbnorm-CT consistently localizes the queried abnormalities and produces more precise, clinically aligned reports.
In Fig. 6, we compare OmniAbnorm-CT with LLaVA-Med and QWen2.5-VL-7B on visual prompted report generation. The baselines show some fundamental mistakes, e.g., LLaVA-Med mislabels the kidney as the gallbladder in the first case, and QWen2.5-VL-7B confuses the liver and gallbladder in the last case. In contrast, OmniAbnorm-CT accurately identifies the marked abnormalities and produces higher-quality, clinically consistent reports.
Appendix B Ablation Studies
We conduct a series of ablation studies on the key factors of OmniAbnorm-CT, from the segmentation module (Section B.1), the context range of input image (Section B.2), to the form of visual prompts (Section B.3). This section introduces the settings and results of these experiments sequentially.
B.1 Ablation Study on Segmentation Module
The integration of the segmentation module allows OmniAbnorm-CT to acquire grounding evidence for abnormal findings and generate descriptions accordingly. To validate its necessity, we conduct an ablation study comparing the full model against a variant without the segmentation module (OmniAbnorm-CT w/o Seg), which directly generates findings.
Text-guided grounded report generation. As shown in Table 5, removing the segmentation module significantly degrades performance on most metrics (23/27 across planes). Furthermore, our manual analysis reveals that, without the grounding evidence, the model struggles to accurately detect the queried abnormalities in CT images.
Grounded report generation. Similarly, Table 6 shows that OmniAbnorm-CT w/o Seg consistently underperforms the full model on all metrics, reaffirming that segmentation-driven evidence is essential for accurate grounding and generating clinically faithful reports.
Model BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
@@ B.1 Ablation Study on Segmentation Module
OmniAbnorm-CT w/o Seg 4.50 2.33 1.21 20.53 31.06 7.04 8.50 7.02 7.25
OmniAbnorm-CT 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
@@
OmniAbnorm-CT w/o Seg 4.50 2.47 1.27 19.07 28.34 6.27 7.96 6.43 6.17
OmniAbnorm-CT 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
@@
OmniAbnorm-CT w/o Seg 4.17 2.16 1.12 16.62 25.43 5.84 7.06 5.53 5.67
OmniAbnorm-CT 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
Table 5: Ablation study of the segmentation module for text-guided grounded report generation task.
Model BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
@@
Table 6: Ablation study of the segmentation module in grounded report generation task.
B.2 Ablation Study on Context Range
In clinical practice, adjacent-slice context is important for CT interpretation, including segmentation and diagnosis [nnUNET, zheng2024large]. To assess its impact on report generation, we vary the number of input slices in visual prompted report generation, using a base QWen2.5-VL-7B. Surprisingly, Table 7 shows that adding slices yields no consistent gains across metrics. We hypothesize that current LVLMs, which encode the slices independently into tokens, struggle to model subtle inter-slice spatial relations. Accordingly, we use only the center slice as input to the LVLM in subsequent training and evaluation of OmniAbnorm-CT.
Input Slices BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadGraph